# Estimating the linear relation between variables that are never jointly observed: an application in in vivo experiments

### v1

@@ Abstract
This work is motivated by in vivo experiments in which measurement are destructive so that the variables of interest can never be observed simultaneously when the aim is to estimate the regression coefficients of a linear regression. Assuming that the global experiment can be decomposed into sub experiments (corresponding for example to different doses) with distinct first moments, we propose different estimators of the linear regression which take account of that additional information. We consider estimators based on moments as well as estimators based optimal transport theory. These estimators are proved to be consistent as well as asymptotically Gaussian under weak hypotheses. The asymptotic variance has no explicit expression, except in some particular cases, and specific bootstrap approaches are developed to build confidence intervals for the estimated parameter. A Monte Carlo study is conducted to assess and compare the finite sample performances of the different approaches.
1 Motivation and context
@@
In order to study the effects of a certain treatment on a living organism, in vivo experiments are often conducted. In the context of a complex organism response, scientists may be interested in studying multiple variables describing the effect from different perspectives. In particular, such variables of interest often include a macroscopic biomarker only available through in vivo data on the one hand, and a microscopic biomarker that can also be observed on a cellular level. The interest in this case lies in predicting the former with the latter. For instance, in the context of studying the adverse effects induced by radiotherapy on healthy tissues, the potential outcomes of interest may manifest in the form of lesions, that are quantified in a form a certain macroscopic biomarker, and the levels of a some predictor such as gene expression. These measures often require sacrificing the animal. As a result, the quantities of interest cannot be observed on the same animals. Since the goal is to establish relationships between these variables, a problem of statistical data fusion arises.
In this work we propose an approach to estimate the relationship between the variables that are not simultaneously observed under the experimental setting described above, based on a conditional model assuming linear relationship within every component (experimental condition). An estimator derived with the method of moments as well as optimal transport solution using Wasserstein distance are considered for the real data problem in question.
1.1 Example of an in vivo experiment
@@
Figure 1: Schematic representation of the design of an in vivo experiment studying the effect of irradiated volume.
@@
An illustrative example of an in vivo experiment where the variables of interest are never observed simultaneously is presented in Figure 1. In this experiment, mice are irradiated on the lungs with different volume, with a goal of studying the role of irradiated volume in the appearing of radio-induced adverse effects. The latter are assessed by measuring septal thickening, a macroscopic biomarker of radio-induced adverse effects in the lungs. The other variable that is measured with the purpose of predicting the adverse effect related variable is the expression of multiple pro-inflammatory genes. As shown in Figure 1, there are two independent cohorts in the study, one is used to measure gene expression, whereas the other for measuring septal thickness. This is a results of both measures being of destructive nature, which does not allow them to be taken on the same animals.
@@
Comparing distributions of the measurements arising from the two cohorts, such as those presented in Figure 2, one may suspect a correlation or even a linear relationship between the variables. In order to assert whether such a relationship exists, one has to connect two variables that are never observed simultaneously, which translates into solving a data fusion problem. In this example, there are four groups indicating time points (1, 3, 6 and 12 months after irradiation) when the corresponding animals are sacrificed and the measurements are taken. Thus, the categorical variable indication the time point, which is known for every observation, can be used as an additional variable in order to link the predictor and the predicted variables between each other.
@@
Figure 2: Distribution of the data, collected from the irradiated patch under SBRT with 3 mm beam size: the expression of the gene IL6 on the left, and septal thickness on the right. The measurements were made 1, 3, 6 and 12 months after irradiation.
1.2 Existing research
@@
The task of linking variables that are never jointly observed cannot be approached as a typical missing values problem, since most methods for inference on incomplete data require a sufficient overlap, which is completely absent in the case treated here. As a result, any approaches applying related frameworks such as multiple imputation in the context of data fusion are unsuitable for our application. For example, Carrig et al., (2015) use multiple imputation to integrate disparate datasets allowing for the absence of the overlap, but requiring a calibration dataset, where all the variables of interest must be jointly observed.
@@
Other approaches to data fusion available in literature include factor analysis (Cudeck,, 2000), statistical matching (Mitsuhiro and Hoshino,, 2020), Bayesian network inference (Triantafillou et al.,, 2010; Tsamardinos et al.,, 2012) and Gaussian Markov combinations (Massa and Riccomagno,, 2017). These methods are designed for linking variables that are not observed simultaneously through covariates, present for both variables of interest. This corresponds to the properties of the in vivo data described in Section 1.1. However, the covariates in these approaches are random variables, typically continuous, and often assumed to be Gaussian, which is the case in Cudeck, (2000) and Massa and Riccomagno, (2017). The groups variable available through in vivo experiments may present in a continuous form, but the presence of such categories as control and sham makes it impossible to assume continuity and normality. The Bayesian network approaches introduced by Triantafillou et al., (2010) and Tsamardinos et al., (2012), aimed at inferring binary causal relationships between variables, are more suitable for large datasets with a high number of covariates. Finally, currently available research in statistical matching addresses such aspects as not-at-random missingness (Mitsuhiro and Hoshino,, 2021) and high dimensionality (Mitsuhiro and Hoshino,, 2020). This approach is based on the idea of comparing distances between the covariates from the datasets of interest, which cannot be done by taking the group variable as the covariate. It can be noted that the goal of the aforementioned examples in statistical matching is to group individuals before imputation, which is not necessary in our case since the groups are already known.
2 Estimating the linear relation between variables that are never jointly observed
2.1 Problem and presentation of the different identification approaches
@@
$Y$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}X_{j}+\epsilon.$ (1)
@@
2.1.1 Moment approach
A first and simple approach is based on first moments identification. Taking the conditional expectation, given $G=k$, in (1), we have for $k=1,\ldots,K$,
$\mu_{Y}^{k}$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k},$ (2)
@@ 2.1.1 Moment approach
We denote by $\boldsymbol{\mu}_{1,X}$ the $K\times(d+1)$ design matrix, whose $k$th row is equal to $(1,\boldsymbol{\mu}_{X}^{k\top})$ with $\boldsymbol{\mu}_{X}^{k}=(\mu_{X_{1}}^{k},\cdots,\mu_{X_{d}}^{k})^{\top}$, by $\boldsymbol{\mu}_{Y}$ the $K$ dimensional vector with elements $(\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$. The $K$ linear equations in (2) can we written in a matrix way, $\boldsymbol{\mu}_{Y}=\boldsymbol{\mu}_{1,X}\boldsymbol{\beta}$.
@@
If model (1) holds and assumption $\mathbf{H}_{1}$ is fulfilled, $\boldsymbol{\beta}$ is uniquely identified in terms of the conditional first order moments of $\mathbf{X}$ and $Y$ given $G$,
@@
where $\sigma_{Y}^{2}$ is the variance of $Y$, $\boldsymbol{\Gamma}_{X}$ is the covariance matrix of $\mathbf{X}$, and $\boldsymbol{\beta}_{-0}=(\beta_{1},\ldots,\beta_{d})$.
@@
2.1.2 Optimal transport and minimum Wasserstein distance approach
@@
If we assume that, given $G=k$, $\mathbf{X}$ is a Gaussian random vector and $\epsilon$ is Gaussian, we have, if the linear model (1) holds, that $Y$ is also Gaussian given $G=k$, with expectation $\mu_{Y}^{k}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k}$ and variance $\sigma_{\epsilon}^{2}+\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}$, where $\boldsymbol{\Gamma}_{X}^{k}$ is the variance matrix of $\mathbf{X}$ given $G=k$. Thus, the Wasserstein distance between $D_{\gamma}$, the distribution of $\gamma_{0}+\boldsymbol{\gamma}_{-0}^{\top}\mathbf{X}+\epsilon$, and $D_{Y}$, the distribution of $Y$, is equal to
@@
$\varphi(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k})^{2}+\left(\sigma_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (3)
@@
If model (1) holds and if assumption $\mathbf{H}_{1}$ is fulfilled, $\varphi(\boldsymbol{\gamma},\sigma^{2})$ has its unique minimum at $\boldsymbol{\gamma}=\boldsymbol{\beta}$ and $\sigma^{2}=\sigma_{\epsilon}^{2}$.
2.2 Sampled data and estimators
@@ 2.2.1 Moment estimators
$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{Y}$
@@
$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}$ (4)
with $\mathbf{w}=(w_{1},\ldots,w_{K})$. The first moment estimator considered just before corresponds to equal weights $w_{k}=K^{-1}$.
@@
$\widehat{\sigma}_{\epsilon}^{2,M}$ $=\widehat{\sigma}_{Y}^{2}-(\widehat{\boldsymbol{\beta}}_{-0}^{M})^{\top}\widehat{\boldsymbol{\Gamma}}_{X}\widehat{\boldsymbol{\beta}}_{-0}^{M}.$ (5)
2.2.2 Minimum Wasserstein distance estimators
@@
$\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}+\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (6)
Note that in absence of an a priori information on the probability $\pi_{k}$ of observing group $k$, we can set $\pi_{k}=K^{-1}$. We denote by $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})$ minimizers of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ which are obtained with iterative optimization algorithms based on gradient descent. The algorithm can be initialized with $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}^{2,M})$
2.3 Consistency and asymptotic distribution
@@
If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, and assumption $\mathbf{H}_{1}$ is fulfilled, the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}_{\epsilon}^{2,M})$ defined by (4) and (5) converges in probability to $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ when $n_{\min}$ tends to infinity.
@@
If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain 0, suppose that model (1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, then the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W}_{\epsilon})$ that minimize (6) converges in probability to $(\beta,\sigma^{2}_{\epsilon})$ when $n_{\min}$ tends to infinity.
@@
$0<c\leq\frac{\max(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}{\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}\leq C<+\infty,$ (7)
@@
If model (1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{4})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain $(0,0)$, then, as $n$ tends to infinity,

### v3

@@ Abstract
In modern experimental science there is a commonly encountered problem of estimating the coefficients of a linear regression in the context where the variables of interest can never be observed simultaneously. Assuming that the global experiment can be decomposed into sub-experiments with distinct first moments, we propose two estimators of the linear regression that take this additional information into account. We consider an estimator based on moments, and an estimator based on optimal transport theory. These estimators are proven to be consistent as well as asymptotically Gaussian under weak hypotheses. The asymptotic variance has no explicit expression, except in some particular cases, for which reason a stratified bootstrap approach is developed to build confidence intervals for the estimated parameters, whose consistency is also shown. A simulation study, assessing and comparing the finite sample performances of these estimators, demonstrated the advantages of the bootstrap approach in multiple realistic scenarios. An application to in vivo experiments, conducted in the context of studying radio-induced adverse effects on mice, revealed important relationships between the biomarkers of interest that could not be identified with the considered naive approach.
1 Introduction
@@
In order to study the effects of a certain treatment on a living organism, in vivo experiments are often conducted. In the context of a complex organism response, scientists may be interested in studying multiple variables describing the effect on different scales. In particular, such variables of interest often include a macroscopic biomarker, only available through in vivo data, and a microscopic biomarker that can also be observed on a cellular level (i.e. in vitro). The interest in this case lies in predicting the former with the latter. For instance, in the context of studying the adverse effects induced by radiotherapy on healthy tissues, the potential outcomes of interest manifest in the form of the severity of macroscopic lesions, and predictors such as gene expression. These measures often require sacrificing the animals that they are taken from. As a result, the quantities of interest cannot be observed on the same animals. Since the goal is to establish relationships between these variables, a problem of statistical data fusion arises.
@@
Figure 1.1: Schematic representation of the design of an in vivo experiment studying the effect of irradiated volume.
@@
An illustrative example of an in vivo experiment where the variables of interest are never observed simultaneously is presented in Figure 1.1. In this experiment, mice are irradiated on the lungs with different volume, with a goal of studying the role of irradiated volume in the appearance of radio-induced adverse effects. The latter are assessed by measuring septal thickening, a histological marker of lung injuries. The other variable that is measured with the purpose of predicting the adverse effect-related variable is the expression of multiple pro-inflammatory genes. As shown in Figure 1.1, there are two independent cohorts in the study, one is used to measure gene expression, whereas the other for measuring septal thickness. This is a results of both measures being of destructive nature, which does not allow them to be taken on the same animals.
@@
Comparing distributions of the measurements arising from the two cohorts, such as those presented in Figure 1.2, one may suspect a correlation or even a linear relationship between the variables. In order to assert whether such a relationship exists, one has to connect two variables that are never observed simultaneously, which translates into solving a data fusion problem. In this example, there are four groups indicating time points (1, 3, 6 and 12 months after irradiation) when the corresponding animals are sacrificed and the measurements are taken. Thus, the categorical variable indicating the time point, which is known for every observation, can be used as an additional variable in order to link the predictor and the predicted variables between each other.
@@
Figure 1.2: Distribution of the data, collected from the irradiated patch under SBRT with 3 mm beam size: the expression of the gene IL6 on the left, and septal thickness on the right. The measurements were made 1, 3, 6 and 12 months after irradiation.
@@
The task of linking variables that are never jointly observed cannot be approached as a typical missing values problem, since most methods for inference on incomplete data require a sufficient overlap, which is completely absent in the case treated here. As a result, any approaches applying frameworks such as multiple imputation in the context of data fusion are unsuitable for our application. For example, Carrig et al., (2015) use multiple imputation to integrate disparate datasets allowing for the absence of the overlap, but requiring a calibration dataset, where all the variables of interest must be jointly observed.
@@
Other approaches to data fusion available in literature include factor analysis (Cudeck,, 2000), statistical matching (Mitsuhiro and Hoshino,, 2020), Bayesian network inference (Triantafillou et al.,, 2010; Tsamardinos et al.,, 2012) and Gaussian Markov combinations (Massa and Riccomagno,, 2017). These methods are designed for linking variables that are not observed simultaneously through covariates, present for both variables of interest. This corresponds to the properties of the in vivo data described above. However, the covariates in these approaches are random variables, typically continuous, and often assumed to be Gaussian, which is the case in Cudeck, (2000) and Massa and Riccomagno, (2017). The grouping variable available through in vivo experiments cannot be represented in a continuous form since categories such as control and sham make it impossible to assume continuity and normality. The Bayesian network approaches introduced by Triantafillou et al., (2010) and Tsamardinos et al., (2012), aimed at inferring binary causal relationships between variables, are more suitable for large datasets with a high number of covariates. Finally, currently available research in statistical matching addresses such aspects as not-at-random missingness (Mitsuhiro and Hoshino,, 2021) and high dimensionality (Mitsuhiro and Hoshino,, 2020). This approach is based on the idea of comparing distances between the covariates from the datasets of interest, which cannot be done by taking the group variable as the covariate. It can be noted that the goal of the aforementioned examples in statistical matching is to group individuals before imputation, which is not necessary in our case since the groups are already known.
In this work we propose an approach to estimate the relationship between the variables that are not simultaneously observed under the experimental setting described above, based on a conditional model assuming linear relationship within every component, defined by the grouping variable that is assumed to be present. An estimator derived with the method of moments as well as optimal transport solution using Wasserstein distance are considered for the real data problem in question. Both approaches do not require any overlap between the two cohorts of the experiment, and are based on weak assumptions ensuring model identifiability and moments existence. To construct confidence intervals, we propose both asymptotic and bootstrap-based results, the former being often unfeasible in practice, with the latter demonstrating practical advantages in numerous realistic settings.
2 Identification approaches
@@
$Y$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}X_{j}+\epsilon.$ (2.1)
@@
2.1 Moment approach
A first and simple approach is based on first moments identification. Taking the conditional expectation, given $G=k$, in (2.1), we have for $k=1,\ldots,K$,
$\mu_{Y}^{k}$ $=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k},$ (2.2)
@@ 2.1 Moment approach
We denote by $\boldsymbol{\mu}_{1,X}$ the $K\times(d+1)$ design matrix, whose $k$th row is equal to $(1,\boldsymbol{\mu}_{X}^{k\top})$ with $\boldsymbol{\mu}_{X}^{k}=(\mu_{X_{1}}^{k},\cdots,\mu_{X_{d}}^{k})^{\top}$, by $\boldsymbol{\mu}_{Y}$ the $K$ dimensional vector with elements $(\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$. The $K$ linear equations in (2.2) can we written in a matrix way, $\boldsymbol{\mu}_{Y}=\boldsymbol{\mu}_{1,X}\boldsymbol{\beta}$.
@@
If model (2.1) holds and assumption $\mathbf{H}_{1}$ is fulfilled, $\boldsymbol{\beta}$ is uniquely identified in terms of the conditional first order moments of $\mathbf{X}$ and $Y$ given $G$,
@@
where $\sigma_{Y}^{2}$ is the variance of $Y$, $\boldsymbol{\Gamma}_{X}$ is the covariance matrix of $\mathbf{X}$ with elements $\mbox{Cov}(X_{i},X_{j})=\sum_{k=1}^{K}\mbox{Cov}(X_{i},X_{j}|G=k)\mathbb{P}[G=k]$ for $i$ and $j$ in $\{1,\cdots,d\}$, and $\boldsymbol{\beta}_{-0}=(\beta_{1},\ldots,\beta_{d})$.
@@
2.2 Optimal transport approach
@@
If we assume that, given $G=k$, $\mathbf{X}$ is a Gaussian random vector and $\epsilon$ is Gaussian, we have, if the linear model (2.1) holds, that $Y$ is also Gaussian given $G=k$, with expectation $\mu_{Y}^{k}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\mu_{X_{j}}^{k}$ and variance $\sigma_{\epsilon}^{2}+\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}$, where $\boldsymbol{\Gamma}_{X}^{k}$ is the variance matrix of $\mathbf{X}$ given $G=k$. Thus, the Wasserstein distance between $D_{\gamma}$, the distribution of $\gamma_{0}+\boldsymbol{\gamma}_{-0}^{\top}\mathbf{X}+\epsilon$, and $D_{Y}$, the distribution of $Y$, is equal to
@@
$\varphi(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k})^{2}+\left(\sigma_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (2.3)
@@
If model (2.1) holds and if assumption $\mathbf{H}_{1}$ is fulfilled, $\varphi(\boldsymbol{\gamma},\sigma^{2})$ has its unique minimum at $\boldsymbol{\gamma}=\boldsymbol{\beta}$ and $\sigma^{2}=\sigma_{\epsilon}^{2}$.
3 Sampled data and estimators
@@ 3.1 Moment estimators
$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{Y}$ (3.1)
@@
$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}$ (3.2)
with $\mathbf{w}=(w_{1},\ldots,w_{K})$. The first moment estimator considered in (3.1) corresponds to equal weights $w_{k}=K^{-1}$.
@@
$\widehat{\sigma}_{\epsilon}^{2,M}$ $=\widehat{\sigma}_{Y}^{2}-(\widehat{\boldsymbol{\beta}}_{-0}^{M})^{\top}\widehat{\boldsymbol{\Gamma}}_{X}\widehat{\boldsymbol{\beta}}_{-0}^{M}.$ (3.3)
3.2 Optimal transport estimators
@@
$\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}+\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (3.4)
Note that in absence of a priori information on the probability $\pi_{k}$ of observing group $k$, we can set $\pi_{k}=K^{-1}$. We denote by $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})$ minimizers of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ which are obtained with iterative optimization algorithms based on gradient descent. The algorithm can be initialized with $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}^{2,M})$
4 Consistency and asymptotic distribution
@@
If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, and assumption $\mathbf{H}_{1}$ is fulfilled, the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}_{\epsilon}^{2,M})$ defined by (3.2) and (3.3) converges in probability to $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ when $n_{\min}$ tends to infinity.
@@
If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain 0, suppose that model (2.1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, then the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W}_{\epsilon})$ that minimize (3.4) converges in probability to $(\beta,\sigma^{2}_{\epsilon})$ when $n_{\min}$ tends to infinity.
@@
$0<c\leq\frac{\max(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}{\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}\leq C<+\infty,$ (4.1)
@@
If model (2.1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{4})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain $(0,0)$, then, as $n$ tends to infinity,
@@
5 The particular case of simple linear regression
To show the difficulty, let us examine the case of simple linear regression, that is to say $d=1$. The following linear model
$Y$ $=\beta_{0}+\beta_{1}X+\epsilon$
is supposed to hold and if there are two groups $k$ and $j$ such that $\mu_{X}^{k}\neq\mu_{X}^{j}$, the identification assumption $\mathbf{H}_{1}$ is fulfilled and $\beta_{0}$ and $\beta_{1}$ can be uniquely determined. The moment estimators of $\beta_{0}$ and $\beta_{1}$ defined in (3.2) have simple expressions,
$\widehat{\beta}_{0}$ $=\widehat{\mu}_{Y,w}-\widehat{\beta}_{1}\widehat{\mu}_{X,w}$ (5.1)
$\widehat{\beta}_{1}$ $=\frac{\sum_{k=1}^{K}w_{k}\widehat{\mu}_{X}^{k}\widehat{\mu}_{Y}^{k}-\widehat{\mu}_{X,w}\widehat{\mu}_{Y,w}}{\sum_{k=1}^{K}w_{k}\left(\widehat{\mu}_{X}^{k}\right)^{2}-\widehat{\mu}_{X,w}^{2}}$ (5.2)
[…]
We focus on the asymptotic variance of the estimator $\widehat{\beta}_{1}$ of the slope parameter $\beta_{1}$, which is often the parameter of interest. We have
Lemma 5.1.
Suppose that the model (2.1) is true, with $d=1$ and $K\geq 2$. If there are two groups $k$ and $j$ such that $\mu_{X}^{k}\neq\mu_{X}^{j}$ the vector $(\beta_{0},\beta_{1})$ is identifiable and, as $n$ tends to infinity,
$\sqrt{n}\left(\widehat{\beta}_{1}-\beta_{1}\right)$ $\rightsquigarrow\mathcal{N}(0,\sigma^{2}_{\beta_{1}})$
with
$\sigma^{2}_{\beta_{1}}$ $=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left(\beta_{1}^{2}\sigma^{2}_{X,k}+\sigma^{2}_{Y,k}\right)\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}$
and $\mu_{X,w}=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}$ and $\mbox{Var}_{w}(X)=\sum_{k=1}^{K}w_{k}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}$.
Lemma 5.1 shows us that even in a very simple framework (only one covariate), the asymptotic variance of the estimator of the slope $\beta_{1}$ is quite complicated. It also reveals that minimizing the asymptotic variance $\widehat{\beta}_{1}$ with respect to the weights $w_{k}$ is not a simple task.